Presented below are condensed financial statements adapted from those of two actual companies competing in the pharmaceutical industry—Johnson and Johnson (J&J) and Pfizer, Inc. ($ in millions, except per share amounts).

Balance Sheets
($ in millions, except per share data)
J&J Pfizer
Assets:
Cash $ 14,733 $ 9,453
Short-term investments 5,775 12,031
Accounts receivable (net) 8,459 10,660
Inventories 5,336 8,865
Other current assets 5,325 5,190
Current assets 39,628 46,199
Property, plant, and equipment (net) 14,474 22,915
Intangibles and other assets 18,750 72,075
Total assets $ 72,852 $ 141,189
Liabilities and Shareholders' Equity:
Accounts payable $ 6,721 $ 4,356
Short-term notes 4,142 11,821
Other current liabilities 8,968 13,863
Current liabilities 19,831 30,040
Long-term debt 4,580 7,380
Other long-term liabilities 6,590 23,585
Total liabilities 31,001 61,005
Capital stock (par and additional paid-in capital) 5,070 69,000
Retained earnings 45,368 43,747
Accumulated other comprehensive income (loss) (850 ) 260
Less: Treasury stock and other equity adjustments (7,737 ) (32,823 )
Total shareholders' equity 41,851 80,184
Total liabilities and shareholders' equity $ 72,852 $ 141,189
Income Statements
Net sales $ 48,817 $ 52,143
Cost of goods sold 13,814 11,470
Gross profit 35,003 40,673
Operating expenses 21,349 30,072
Other (income) expense—net (580 ) 3,805
Income before taxes 14,234 6,796
Tax expense 4,270 2,039
Net income $ 9,964 $ 4,757 *
Basic net income per share $ 3.07 $ 0.23

* This is before income from discontinued operations.

Evaluate and compare the two companies by responding to the following questions.

Note: Because two-year comparative statements are not provided, you should use year-end balances in place of average balances as appropriate.

Required:
1. For both companies, compute the ratios below.
2. Evaluate and compare the two companies.
Answer #1

I HOPE ITS HELPFUL TO YOU IF YOU HAVE ANY DOUBTS PLS COMMENTS BELOW..I WILL BE THERE TO HELP YOU...ALL THE BEST...!!

AS FOR GIVEN DATA...

1. Receivables Turnover = Net credit Sales / Average Accounts receivables

J&J = 44002 / 7154 = 6.15 times Pfizer = 47328 / 9355 = 5.06 times

2. Average Collection Period = 365 / Receivable turnover ratio

J&J = 365 / 6.15 = 59.35 days Pfizer = 365 / 5.06 = 72.13 days

3. Inventory Turnover = Cost of goods sold / Average Inventory

J&J = 12680 / 4112 = 3.08 times Pfizer = 10336 / 6751 = 1.53 times

4. Average Days in Inventory = 365 / Inventory turnover ratio

J&J = 365 / 3.08 = 118.51 days Pfizer = 365 / 1.53 = 238.56 days

5. Profit Margin = Net Income / Net Sales

J&J = 8061 / 44002 = 18.32% Pfizer = 3044 / 47328 = 6.43%

6. Asset Turnover = Net Sales / Total Assets

J&J = 44002 / 55775 = 78.89 times Pfizer = 47328 / 124287 = 38.08 times

7. Return on Assets = Net Income / Total Assets

J&J = 8061 / 55775 = 14.45% Pfizer = 3044 / 124287 = 2.45%

8. Equity Multiplier = Total Assets / Shareholders Equity

J&J = 55775 / 31425 = 1.77 Pfizer = 124287 / 69933 = 1.78

9. Return on Shareholders' Equity = Net Income / Shareholders Equity

J&J = 8061 / 31425 = 25.65% Pfizer = 3044 / 69933 = 4.35%

Which of the two companies appears more efficient in collecting its accounts receivable and managing its inventory? - J&J is efficient in collecting account receivable as its collection period is lesser than Pfizer (59 vs 72 days). J&J is efficient in manageing inventory as well (118 days vs 238 days)

Which of the two firms had greater earnings relative to resources available? - J&J has greater earnings with a ROA of 14.45%

From the perspective of a common shareholder, which of the two firms provided a greater rate of return? - J&J has higher return on Shareholders equity. We can also refer to basic net income per share in the Income statement.

From the perspective of a common shareholder, which of the two firms appears to be using leverage more effectively to provide a return to shareholders above the rate of return on assets? - Leverage is almost same for both i.e. equity multiplier but J&J is earning higher return than Pfizer, so it is using leverage more effectively.